OpenStudy (anonymous):

I'm supposed to solve for all radian solutions, but I have no idea where to begin with this one. I'd really appreciate a push in the right direction. sin(x) + 2sin(x)cos(x) = 0

OpenStudy (anonymous):

sin(x) can be factored out

OpenStudy (anonymous):

if you factor out sin(x), what do you get?

OpenStudy (anonymous):

sin(x) * (1 + cos(x))

OpenStudy (anonymous):

1+2cos(x) @Laine

OpenStudy (anonymous):

you forgot the 2

OpenStudy (anonymous):

oops forgot the 2 :)

OpenStudy (anonymous):

now thats equal to 0 that means sin(x)=0 and 1+2cos(x)=0

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\(\sin x+2\sin x\cos x=0\\(\sin x)(1+2\cos x)=0\)so either \(\sin x=0\) or \(1+2\cos x=0\) for our equation to hold. Solve both (I solved in the interval \((0,2\pi)\)):$$\sin x=0\\x=\pi$$and$$1+2\cos x=0\\2\cos x=-1\\\cos x=-\frac12\\x\in\left\{\frac{2\pi}3,\frac{4\pi}3\right\}$$

OpenStudy (anonymous):

oops, I meant \(x\in\{0,\pi\}\) for the solution of \(\sin x=0\).

OpenStudy (anonymous):

Thank you very much.